OpenStudy (anonymous):

Find the volume of the region R enclosed by the curves y=3sqrt(sin2x) y=0 and the rotated about the x-axis?

OpenStudy (perl):

is this the curve $$\Large y = 3 \sqrt{\sin(2x)}$$

OpenStudy (anonymous):

yes

OpenStudy (perl):

that is not enough conditions to bound the region https://www.desmos.com/calculator/kyfufafhjd

OpenStudy (perl):

can you take a screenshot of the question

OpenStudy (perl):

the region is infinite. you can't find the volume of an infinite region

OpenStudy (anonymous):

Ok, one minute

OpenStudy (perl):

yes i see, hmm. but do you see my point about the region, https://www.desmos.com/calculator/9vnesnm7ar The region is endless

OpenStudy (perl):

http://prntscr.com/73gr91

OpenStudy (anonymous):

yeah I do understand, I had no problem in using the formula. I had the problem in finding the boundaries

OpenStudy (anonymous):

But anyway, thanks for you help!!!!

OpenStudy (perl):

what does it say after 'and the '

OpenStudy (perl):

"Find the volume of the region R enclosed by the curves y = 3 sqrt(sin(2x)) y = 0 and the < missing words> rotated about the x axis ."

OpenStudy (anonymous):

that's all it says, theres nothing in the middle

OpenStudy (perl):

that doesn't make sense grammatically

OpenStudy (anonymous):

yeah I know, but that's how my professor wrote it out lol

OpenStudy (perl):

if you want to salvage the problem, you can assume that the region is bounded by x =0 and x = pi/2. that seems like a good choice. then do $$ \Large { \int_{0}^{\pi / 2 } \pi ~ \left(3\sqrt{\sin(2x)} \right)^2~ dx } $$

OpenStudy (amistre64):

gabriels trumpet suggests that you can find the volume of an infinite region.

OpenStudy (amistre64):

\[\int_{1}^{\infty}x^{-1}~dx=ln(\infty)-\ln(1)=\infty \] \[\int_{1}^{\infty}\pi (x^{-1})^2~dx=-\pi(\infty^{-1}-1^{-1})=\pi \]

OpenStudy (amistre64):

as such, if we have a can of paint of volume pi, we can fill the volume of the rotation, but we dont have enough paint to do a cross section of it :)